Ahmad has enough money to buy 24 apples. If the price of each apple is reduced by 5 cents, he will be able to buy an extra 6 apples with the same amount of money.Find the original cost of each apple.
Found 3 solutions by 092/093, kiru_khandelwal, ntnk:
Answer by 092/093(1)
Answer by kiru_khandelwal(79)

Let 88 be 22% of x
Now 22% of x = (22/100)*x
=> 88 = (22/100)*x
=> 88 = 22x/100
=> 8800 = 22x
=> x = 8800/22 = 400
So 88 is 22% of 400...:-)

Answer by ntnk(50)

88 is 22% of the number 400. This can be found out by dividing 88 by .22 to get 400.

Remember to check: .22 times 400 equals 88.

Solved By: NtNk